the ticket for a dinner/dance are sold for $55 each.
if the dinner costs $13.75 per head, music costs $540,
and the hall rental costs $559, how many ticket must be
sold to have a minimum profit of $5,425?

Answer by mananth(16946)

number of tickets sold be x
..
55x is the revenue
..
dinner = 13.75x
music =540
hall 559
..
55x-13.75x-540-559=5425
41.25x= 5425+540+559
41.25x= 6524
/41.25
x= 158.1
159 tickets must be sold